OpenStudy (anonymous):

Solving radical equations. 8√y=2

OpenStudy (pigrenex):

y = 1/16

OpenStudy (pigrenex):

\[\sqrt{1/16} = 1/4\] so \[8 \times 1/4 = 2\]